# Constraint-Aware Zero-Shot Vision-Language Navigation in Continuous Environments

### v1

@@ IV-D Quantitative and Qualitative Analysis
Impact of Instruction Complexity. To analyze the impact of instruction complexity on navigation performance we divide the instructions of R2R val-unseen split into categories based on the number of sub-instructions, ranging from 1 to 7. For each category, we record the number of success and failure trajectories and calculated the corresponding success rate (SR). The results in Figure 6 reveal a trend that as the number of sub-instructions increases, the success rate (SR) generally decreases. As the number of sub-instructions increases, the system is required to perform more actions and make more decisions, which increases the chance for errors in task execution. This becomes particularly apparent in trajectories with 5 or more sub-instructions, where the SR starts to decrease significantly. This indicates that instruction switching may not always be accurate as the perception model is imperfect. When the number of sub-instructions increases errors tend to accumulate. Enhancing the CSM further could be a promising direction for improvement.
@@
Navigation Visualization. As shown in Figure 7 and Figure 9, we visualize the navigation process of CA-Nav, FBE-Nav, and NavGPT-CE on the same episode. Note that the FBE-Nav uses frontier-based exploration for waypoint selection, while the rest of the process is the same as the CA-Nav. Among these three methods, CA-Nav shows the most stable and coherent navigation, thanks to CSM’s effective instruction decomposition and constraint identification, along with CVM’s comprehensive use of the value map.
@@
Both OpenNav and NavGPT-CE are zero-shot methods that use a waypoint predictor to discretize continuous environments. They rely on caption models to describe observations and make navigation decisions entirely in the text space, based on LLMs. We analyze NavGPT-CE in detail and identify three main types of errors that contribute to its failures: waypoint predictor failure, LLM hallucination, and caption model inaccuracy. As shown in Figure 9, the agent initially follows the instructions correctly, but errors start to appear from step 30. At step 30, the ground truth action was to move forward, and the LLM correctly identified this. However, the waypoint predictor failed to identify the navigable viewpoint, causing the agent to move to the right instead. At step 41, the LLM exhibits context inconsistency hallucination in summarizing observation descriptions. Specifically, the dining room is to the left, which is the correct direction, but the LLM incorrectly summarized it as being in front of the robot. Moreover, since the dining room to the left is too far away, the waypoint predictor failed to identify valid waypoints, leading the LLM to choose to go to the right instead. At step 55, the caption model describes the current panoramic observation as “a bathroom with a toilet and a sink.” However, the primary objects in the scene are stairs, and this incorrect caption ultimately misleads the agent. Our analysis reveals that NavGPT-CE relies solely on converting visual observations into text-based scene descriptions and historical summaries for navigation decisions, making it prone to errors when these descriptions are inaccurate.
@@ IV-E Real Robot Experiments
We conduct real robot experiments based on the QiZhi mobile robot, as shown in Figure 8 (e).
Figure 8: Real robot and real-world scenes.
We equip the robot with a laptop (including an Intel i9-14900HX CPU and a GeForce RTX 4090 GPU) and a Kinect V2.0 RGB-D camera whose HFoV is 84 and VFoV is 42. However, due to the camera’s limited depth sensing range and the inaccuracies or depth loss at the edges, we utilize Depth Anything V2 [82] to generate depth images. An RPLIDAR-A2M8 LiDAR is also used to obtain a relatively accurate pose through Hector SLAM. It is important to note that Hector SLAM is employed solely for estimating the camera’s pose, not for constructing a pre-built map. The robot has a radius of 22.5cm and a height of 137cm.
@@
Figure 9: Navigation visualization of NavGPT-CE for the same episode as shown in Figure 7.
@@
To evaluate the effectiveness of CA-Nav we conduct experiments in diverse indoor scenes including lounge, meeting room, reception room, entrance hall, and office, as shown in Figure 8. These diverse environments test the adaptability of our method to different spatial layouts and semantic contexts. For instructions, we design 8 instructions with different complexity. Easy instructions contain only one sub-instruction and the destination is obvious, such as, “Go to the door.”. Complex instructions are longer and with more than three constraints and the agent can not directly see the destination from its initial position, thus requiring exploration following the instruction. To check our method’s ability to novel landmarks we design open vocabulary destinations such as “robot” and “world cup trophy” (details in Table IX).
Figure 10: Visualization of the real-world navigation
TABLE IX: Information about real robot experiments. The robot’s initial pose will be different in each episode. $d$ represents the distances between the initial position and the destination. SR is the success rate.
Instruction Example Pose $d$ SR
Go to the door. [Uncaptioned image] 5.4m 8/10
Turn left then walk towards the signboard and wait by the elevator. [Uncaptioned image] 5.5m 4/10
Walk out of the door then stop in front of the plant. [Uncaptioned image] 6.6m 5/10
Walk out to the corridor, and stop in front of the poster. [Uncaptioned image] 4.3m 6/10
Walk towards the robot. [Uncaptioned image] 3.6m 7/10
Walk towards the living room then stop beside the couch [Uncaptioned image] 6.0m 4/10
Turn slightly right and walk into the meeting room, step forward then stop in front of the table. [Uncaptioned image] 3.8m 4/10
Go towards the plant, turn right, walk along the wall then stop near the World Cup Trophy. [Uncaptioned image] 10.4m 2/10
@@
For each instruction, we run 10 episodes, with the robot’s initial pose different each time (details in Table IX). The results indicate that even without a pre-built map, our real-time constructed constraint-aware value map is capable of handling long-horizon navigation tasks. Furthermore, thanks to the VLM, CA-Nav demonstrates a certain level of generalization to open vocabularies. As shown in Figure 10, the agent follows a complex instruction consisting of four sub-instructions and ultimately stops successfully near a World Cup Trophy. This indicates that the CA-Nav can generalize to new instructions and effectively track the navigation process.

### v2

@@ IV-D Quantitative and Qualitative Analysis
Impact of Instruction Complexity. To analyze the impact of instruction complexity on navigation performance we divide the instructions of R2R-CE val-unseen split into categories based on the number of sub-instructions, ranging from 1 to 7. For each category, we record the number of success and failure trajectories and calculated the corresponding success rate (SR). The results in Figure 6 reveal a trend that as the number of sub-instructions increases, the success rate (SR) generally decreases. As the number of sub-instructions increases, the system is required to perform more actions and make more decisions, which increases the chance for errors in task execution. This becomes particularly apparent in trajectories with 5 or more sub-instructions, where the SR starts to decrease significantly.
@@
This indicates that instruction switching may not always be accurate as the perception model is imperfect. When the number of sub-instructions increases errors tend to accumulate. Enhancing the CSM further could be a promising direction for improvement.
@@
Constraint Fulfillment Evaluation. To demonstrate the effectiveness of the CSM module, we evaluate the fulfillment of the current constraints through human evaluation. We select the top 25 trajectories from the successful set based on SPL in descending order to assess whether the episodes that best follow the instructions also achieve the highest constraint fulfillment. Additionally, we randomly sample 25 trajectories from the oracle successful set to examine whether constraint fulfillment declines in episodes that reach the ground truth destination but fail to stop correctly. For failure trajectories, to ensure a diverse coverage of different trajectory lengths, we randomly select 25 samples each for short, medium, and long trajectories. Then we calculate the Constraint Fulfillment Score (i.e., CFS) of episodes and CFS for the three type of constraints (D: Direction, O: Object, L: Location) following equation 8. Specifically, $M$ represents the number of the evaluation set which is 25, $N_{j}$ is the number of constraints in the $j$th episode and $C_{j}$ represents the number of specific constraints of type direction, object or location in the $j$th episode. We manually inspected each step of these trajectories and their constraint fulfillment status. A step was considered fulfilled if and only if the CSM determined the constraint was met and manual verification confirmed its correctness; otherwise, it was labeled as a violation.
$CFS_{\text{episode}}$ $=\frac{1}{M}\sum^{M}_{j=1}\frac{1}{N_{j}}\sum^{N_{j}}_{i=1}\mathbb{I}(\text{constraint}_{i})$ (8)
$CFS_{D/O/L}$ $=\frac{1}{M}\sum^{M}_{j=1}\frac{1}{C_{j}}\sum^{C_{j}}_{i=1}\mathbb{I}(\text{constraint}_{i})$
The results are shown in Table IV. Successful trajectories achieve the highest episode-level CFS (0.69) and lead in CFS across all constraint types, demonstrating the effectiveness of CSM in adhering to constraints. In contrast, Oracle Success trajectories exhibit a lower $\text{CFS}_{episode}$ (0.50) and reduced CFS for individual constraint types, suggesting that while the agent reaches the correct location, it follows constraints less rigorously. As failure severity increases, both $\text{CFS}_{episode}$ and $\text{CFS}_{D/O/L}$ drop below 0.5, indicating that CSM fails most of the time. Overall, these results highlight the relationship between CSM’s effectiveness and navigation performance. Episodes with higher navigation success rates tend to exhibit greater constraint fulfillment, and this positive correlation to some extent indicates the effectiveness of CSM.
TABLE IV: Constraint Fulfillment Evaluation.
Type $\text{CFS}_{\text{episode}}$uparrow $\text{CFS}_{D}$uparrow $\text{CFS}_{O}$uparrow $\text{CFS}_{L}$uparrow
Success 0.69 1.00 0.77 0.52
Oracle Success 0.50 0.43 0.50 0.55
Fail(short) 0.42 0.61 0.42 0.23
Fail(medium) 0.38 0.40 0.42 0.36
Fail(long) 0.34 0.31 0.39 0.31
Figure 8: Failure cases visualization. The red font indicates the sub-instruction currently being executed.
Navigation Visualization. As shown in Figure 7, and Figure 1 in Appendix, we visualize the navigation process of CA-Nav, FBE-Nav and NavGPT-CE on the same episode. Note that the FBE-Nav uses frontier-based exploration for waypoint selection, while the rest of the process is the same as the CA-Nav. Among these three methods, CA-Nav shows the most stable and coherent navigation, thanks to CSM’s effective instruction decomposition and constraint identification, along with CVM’s comprehensive use of the value map.
@@
Failure Analysis. To gain deeper insights into the limitations of our approach, we conduct a detailed failure analysis and categorize the four primary failure modes, as illustrated in Figure 8. First, visual reconstruction errors in the simulation environment often lead to failures in object detection. Inaccuracies in scene rendering may distort object appearances, causing the detection model to miss or misinterpret key objects. Second, object detection model errors can arise due to semantic ambiguities. Models such as GroundingDINO may misclassify objects with similar visual and contextual features, leading to incorrect navigation decisions. For example, the bar is mistaken for the message table. Third, location recognition failures occur when vision-language models such as BLIP2-VQA provide incorrect scene interpretations. Despite the presence of clear visual cues, the model may fail to recognize specific locations. For instance, BLIP2 might not correctly identify a bathroom even when a toilet is visible, resulting in incorrect localization. Lastly, constraint decomposition errors occasionally arise from the large language model (LLM). In some cases, the LLM fails to accurately parse and decompose instructions, leading to omissions of constraints. For example, the instruction includes a step of “turn left at table clock”, however, the LLM overlooked this sub-instruction and failed to extract the direction constraint. Such errors can disrupt navigation planning and lead to deviations from the intended path. These failure cases highlight the challenges of integrating multiple vision and language models for embodied navigation and suggest directions for future improvement, such as enhancing simulator’s visual fidelity, refining spatial reasoning capabilities, and improving LLM’s understanding of instructions.
@@ IV-E Real Robot Experiments
We conduct real robot experiments based on the QiZhi mobile robot, as shown in Figure 9 (e). We equip the robot with a laptop (including an Intel i9-14900HX CPU and a GeForce RTX 4090 GPU) and a Kinect V2.0 RGB-D camera whose HFoV is 84 and VFoV is 42. However, due to the camera’s limited depth sensing range and the inaccuracies or depth loss at the edges, we utilize Depth Anything V2 [83] to generate depth images. An RPLIDAR-A2M8 LiDAR is also used to obtain a relatively accurate pose through Hector SLAM. It is important to note that Hector SLAM is employed solely for estimating the camera’s pose, not for constructing a pre-built map. The robot has a radius of 22.5cm and a height of 137cm.
@@
Figure 9: Real-world scenes and real robot.
@@
To evaluate the effectiveness of CA-Nav we conduct experiments in diverse indoor scenes including lounge, meeting room, reception room, entrance hall, and office, as shown in Figure 9. These diverse environments test the adaptability of our method to different spatial layouts and semantic contexts. For instructions, we design 8 instructions with different complexity. Easy instructions contain only one sub-instruction and the destination is obvious, such as, “Go to the door.”. Complex instructions are longer and with more than three constraints and the agent can not directly see the destination from its initial position, thus requiring exploration following the instruction. To check our method’s ability to novel landmarks we design open vocabulary destinations such as “robot” and “world cup trophy” (details in Table V).
@@
For each instruction, we run 10 episodes, with the robot’s initial pose different each time (details in Table V). We deploy CoW-OWL as the baseline. For CA-Nav, we use Depth Anything V2 as the default setting and also compare it with ZoeDepth [85]. The results indicate that CoW can handle simple cases where the target object is directly visible from the initial pose and the instruction consists of only a single sub-step. However, when the instruction requires sequential exploration of multiple objects, CoW fails in most cases. In contrast, CA-Nav effectively handles long-horizon navigation tasks even without a pre-built map, demonstrating the effectiveness of our real-time constructed constraint-aware value map. Furthermore, Depth Anything V2 and ZoeDepth exhibit comparable performance. However, Depth Anything V2 holds a slight advantage, likely due to its superior depth estimation capability. In Appendix, we provide visualizations of CoW and ZoeDepth along with corresponding discussions. Furthermore, thanks to the VLM, CA-Nav demonstrates a certain level of generalization to open vocabularies. As shown in Figure 2 in Appendix, the agent follows a complex instruction consisting of four sub-instructions and ultimately stops successfully near a World Cup Trophy. This indicates that CA-Nav can generalize to new instructions and adapt to novel environments while effectively tracking the navigation process.
TABLE V: Information about real robot experiments. The robot’s initial pose will be different in each episode. $d$ represents the distance between the initial position and the destination. SR is the success rate. CoW represents the detector based version (i.e., CoW-OWL), Ours represents CA-Nav using Depth Anything V2, ZoeDepth represents CA-Nav using ZoeDepth.
Instruction Go to the door. Turn left then walk towards the signboard and wait by the elevator. Walk out of the door then stop in front of the plant. Walk out to the corridor, and stop in front of the poster.
Example Pose [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Distance ($d$) 5.4m 5.5m 6.6m 4.3m
SR: CoW, Ours, ZoeD 4/10, 8/10, 8/10 0/10, 4/10, 3/10 1/10, 5/10, 2/10 2/10, 6/10, 4/10
Instruction Walk towards the robot. Walk towards the living room then stop beside the couch. Turn slightly right and walk into the meeting room, step forward then stop in front of the table. Go towards the plant, turn right, walk along the wall then stop near the World Cup Trophy.
Example Pose [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Distance ($d$) 3.6m 6.0m 3.8m 10.4m
SR: CoW, Ours, ZoeD 6/10, 7/10, 6/10 1/10, 4/10, 3/10 0/10, 4/10, 4/10 0/10, 2/10, 1/10